# Classification of superpotentials for cohomogeneity one Ricci solitons

### v1

@@ 3.1. Classification of superpotentials
together implies that $v_{1}=\pm 1$, $v_{2}=\pm 1$, $d_{1}=d_{2}=2$, and also that $\textbf{v}=\pm(1^{1},-1^{2})$. This implies that all vertices of $\mathrm{conv}(\mathcal{W})$ lie in the two planes
@@
Thus, there can only be type III vertices in $\mathrm{conv}(\mathcal{W})$. In fact, there are no type II vectors in $\mathcal{W}$. To see this, assume for sake of contradiction that $(1^{1},-1^{2},-1^{3})\in\mathcal{W}$. Then by Proposition 4.1(i) and (vii) of [4], $(1^{1},-2^{2}),(-2^{1},1^{2}),(1^{1},-2^{3}),(-2^{1},1^{3}),(1^{2},-2^{3}),(-2^{2},1^{3})\in\mathcal{W}$. As a result, $d_{1}=d_{2}=d_{3}=4$, and (3.4),(3.5) require that for each $i,j=1,2,3$, $i\neq j$,
@@
On the other side, we first suppose that $(1^{1},-1^{2},-1^{i})$ is a vertex. Then $d_{1}=2$, in which case all vertices $\textbf{w}^{\prime}$ of $\mathrm{conv}(\mathcal{W})$, with the possible exception of $\textbf{w}_{0}$, must satisfy $\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}=1$. However, this is impossible, as $(-1^{2})\in\{\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}<1\}$ but $\textbf{w}_{0}\in\{\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}>1\}$. The only remaining possibility is that $(1^{1},-2^{i})$ is a vertex, so that $d_{1}=1$ by (3.8). Similarly, $(-1^{2})\in\{w^{\prime}_{1}-\frac{w^{\prime}_{2}}{2}>1\}$ and $(1^{1},-2^{2})\in\{w^{\prime}_{1}-\frac{w^{\prime}_{2}}{2}<1\}$, which is a contradiction. Therefore there are no type II vectors with nonzero entries in the first two positions in $\mathcal{W}$.
Consider the plane spanned by the first 3 coordinates (see Figure 1). There are no type II vectors, and furthermore (3.8) implies that $(-2^{1},1^{3}),(-2^{1},1^{2})\notin\mathcal{W}$. Since $(-1^{1})$ cannot be a vertex, we have $d_{1}=1$. Let $\textbf{c}_{\pm}=\frac{1}{2}(\textbf{d}+\textbf{v}_{\pm})$. Since $\textbf{v}_{+}+\textbf{v}_{-}=2\textbf{w}_{0}=(2^{1},-4^{2})$, we must in particular have $v_{+,1}+v_{-,1}=2$, but also since both $\frac{1}{2}(\textbf{d}+\textbf{v}_{\pm})$ are null and $d_{1}=1$, we must have $v_{\pm,1}^{2}<1$, which is a contradiction. Thus $(a,b)=(1,1)$ is impossible.
@@
If $(a,b)=(1,0)$, then the only set of weight vectors admitting a superpotential is the Bérard-Bergery–Calabi ansatz.
@@
Let $\textbf{c}=\frac{1}{2}(\textbf{d}+\textbf{w})$ be a (non-null) vertex which is adjacent to $\frac{1}{2}(\textbf{d}+\textbf{v})$. It is also adjacent to $\frac{1}{2}(\textbf{d}-\textbf{v})$, giving us that $J(\textbf{d}-\textbf{v},\textbf{d}+\textbf{w})=0$. This implies that $J(\textbf{d}+\textbf{v},\textbf{d}+\textbf{w})\neq 0$, so either $\textbf{c},\frac{1}{2}(\textbf{d}+\textbf{v})$ do not satisfy the unique sum condition, or $\frac{1}{2}(\textbf{v}+\textbf{w})\in\mathcal{W}$.
@@
Since $\mathrm{conv}(\mathcal{W})$ has nonzero $(r-1)$-dimensional measure, there are at least $(r-1)$ vertices in $\{s=-1\}$ adjacent to $\frac{1}{2}(\textbf{d}+\textbf{v})$. Denote them by $\textbf{w}_{1},\dots,\textbf{w}_{r-1}$. For each $i$, we would then have $\textbf{w}^{\prime}_{i}\equiv\frac{1}{2}(\textbf{v}+\textbf{w}_{i})\in\mathcal{W}$. Thus
@@
Thus the $\{\textbf{w}_{i}\}$ and the $\{\textbf{w}^{\prime}_{i}\}$ lie on parallel affine hyperplanes of $\{s=-1\}$.
@@
Then $d_{1}=d_{2}=2$, $v_{1}=-1$, $v_{2}=1$, and the right-hand side inequality is an equality, so $\textbf{v}=(-1^{1},1^{1})$, contradicting $\sum_{i=1}^{r}v_{i}=-1$. Hence, all three of $v_{1},v_{2},v_{3}$ are equal to ±$\pm 1$. In fact, since $d_{1},d_{2},d_{3}\geq 2$, we must have $v_{1}=-1,v_{2}=1,v_{3}=1$, and the null condition inequality is an equality, implying that $\textbf{v}=(-1^{1},1^{2},1^{3})$, contradicting $\sum_{i=1}^{r}v_{i}=-1$.
Hence, by elimination, each $\textbf{w}_{i}$ is a type III vertex. Suppose without loss of generality that $\textbf{w}_{1}=(1^{1},-2^{2})$. Then
@@
Suppose that $r\geq 3$. Consider for now the equality
@@
The only way in which the difference of two type III vectors can have only even entries is that the $(-1)$’s are in the same position. Since there must be $r-1$ such vectors, the only possibility up to permutation is $\textbf{w}_{i}=(1^{1},-2^{i+1})$, $i=1,\dots,r-1$. All the $\textbf{w}_{i}$ lie on the hyperplane $\{x_{1}=1\}\cap\{s=-1\}$. Since $2\textbf{w}^{\prime}_{i}-\textbf{w}_{i}=\textbf{v}$ for each $i$, we must have $(\textbf{w}^{\prime}_{1})_{1}=(\textbf{w}^{\prime}_{2})_{1}=\dots=(\textbf{w}^{\prime}_{r-1})_{1}$.
@@
First, $d_{1}=1$. Then $v_{1}\in\{-1,0\}$. Since $v_{1}$ must differ from $-(\textbf{w}_{i})_{1}=-1$ by an even integer, we have $v_{1}=-1$. Since $\frac{1}{2}(\textbf{d}+\textbf{v})$ is null, we have $\textbf{v}=(-1,0,\dots,0)$ and we have exactly Example 18 of [2], on the Bérard-Bergery–Calabi ansatz.
@@
In particular, we must have $|v_{1}/d_{1}|\leq 1$, with equality if and only if $v_{1}=-d_{1}$ and $\textbf{v}=(-d_{1},0,\dots,0)$. However, we have assumed that $d_{1}>1$, so equality is impossible. Therefore $|v_{1}/d_{1}|<1$ and $v_{k}/2>0$ for each $k$. The only possibility is $v_{k}=1$ for each $k$, with $v_{1}/d_{1}=-1/2$. The midpoint of $v_{1}$ and $1$ must be a component of a vector in $\mathcal{W}$, so in fact $v_{1}\in\{-1,-3,-5\}$. The second equation implies that in fact $v_{1}=-1$, $d_{1}=2$, so that $\sum_{i=1}^{r}v_{i}\geq 0$, which is a contradiction.
In the case $r=2$, there are only three possibilities for $\textbf{w}^{\prime}_{1}$. First, if $\textbf{w}^{\prime}_{1}=(0,-1)$, then $\textbf{v}=(1,0)$ and we again have the Bérard-Bergery–Calabi ansatz. Otherwise, either $\textbf{w}_{1}^{\prime}=(0,-1)$ and $\textbf{v}=(-3,2)$ or $\textbf{w}_{1}^{\prime}=(-2,1)$ and $\textbf{v}=(-5,4)$; in both of these cases, the null condition for $\frac{1}{2}(\textbf{d}+\textbf{v})$ contradicts (3.10). This concludes the classification for the case $(a,b)=(1,0)$.
@@
Let the two (adjacent) vertices in $\{s>-1\}$ be $\frac{1}{2}(\textbf{d}\pm\textbf{v})$ and let $\frac{1}{2}(\textbf{d}+\textbf{u})$ be the vertex in $\{s<-1\}$. Let $Q_{\pm}=\ker J(\textbf{d}\pm\textbf{v},\textbf{d}+\cdot)\cap\{s=-1\}$ and $Q_{u}=\ker J(\textbf{d}+\textbf{u},\textbf{d}+\cdot)\cap\{s=-1\}$. Each is a hyperplane of $\{s=-1\}$, and $Q_{+}\cap Q_{-}=\emptyset$.
@@
We consider the situation case by case, choosing $\textbf{w}_{\pm}$ to be type I, II and III vectors respectively.
@@
Thus, we have obtained a special case of the Bérard-Bergery–Calabi ansatz discussed in [2]. We thus permute the indices $1\leftrightarrow 3$ to match the notation, so that the geometric data is $(d_{1},d_{2},d_{3})=(1,2,2)$ and $\mathcal{W}=\{(0,-1,0),(0,0,-1),(1,-2,0),(1,0,-2)\}$. However, the superpotential vectors differ from those in Lemma 3.15 and [2, Example 18]. They are $\mathcal{C}=\frac{1}{2}(\textbf{d}+(\mathcal{B},0))$, where
@@
Suppose $\textbf{w}_{+}$ is of type II, say $\textbf{w}_{+}=(1^{1},-1^{2},-1^{3})$, and that $\textbf{w}_{-}$ is also of type II and thus lies in $R$. Say $\textbf{w}_{-}=(-1^{1},1^{2},-1^{3})$, then $\textbf{u}=(-2^{3})$ and $\textbf{v}=(2^{1},-2^{2})$. The null condition requires that $d_{3}=4$. The conditions $J(\textbf{d}+\textbf{w}_{+},\textbf{d}+\textbf{w}_{-})=0$ and $J(\textbf{d}\pm\textbf{v},\textbf{d}\pm\textbf{v})=0$ are equivalent and both reduce to $\frac{4}{d_{1}}+\frac{4}{d_{2}}=1$. Consider again the plane spanned by the first 3 coordinates. Since $\overline{\textbf{w}_{+}\textbf{w}_{-}}$ is an edge, we have $(1^{1},-2^{2}),(1^{1},-2^{3})\notin\mathcal{W}$. Furthermore, $(1^{1},-2^{2},0)$, $(0,-2,1)$ and $(-1^{1},-1^{2},1^{3})$ cannot be vertices adjacent to $\textbf{w}_{+}$, since they are not in $\ker J(\textbf{d}+\textbf{w}_{+},\textbf{d}+\cdot)$. This is a contradiction, as can be seen from Figure 1 and the fact that $(-1^{1},-1^{2},1^{3})\in\mathcal{W}$.
@@
Since $(-1^{2})\in\mathcal{W}$ by Lemma 4.1(ii) of [4], and we have eliminated all other possibilities as can be seen in Figure 1, $(-1^{2})$ must be adjacent to $\textbf{w}_{+}$, and therefore $1=\frac{2}{d_{2}}$ and $d_{2}=2$, contradicting (3.14) and completing the proof of the lemma.
We may summarise Lemmas 3.13, 3.14, 3.15 and 3.16 in the following statement.
@@
    $\mathcal{W}=\emptyset$ with $d_{1}=1$. This is the 2-dimensional Bryant soliton, discussed in [2, Example 15]. There is one superpotential with $\mathcal{C}=\{\frac{1}{2}(\textbf{d}+(1,0)),\frac{1}{2}(\textbf{d}+(-1,0))\}$.
@@
    $\mathcal{W}=\{(-1)\}$ with $d_{1}=4$. This is the 5-dimensional Bryant soliton, discussed in [2, Example 15]. One superpotential is admitted, with $\mathcal{C}=\{\frac{1}{2}\textbf{d},\frac{1}{2}(\textbf{d}+(-2,0))\}$.
@@
    $\mathcal{W}=\{(-1,0),(0,-1)\}$ with $d_{1}=d_{2}=2$. This is the warped product of two Einstein 2-manifolds with positive Einstein constant, discussed in [2, Example 16]. We have a superpotential with $\mathcal{C}=\{\frac{1}{2}(\textbf{d}+(-1,-1,0)),\frac{1}{2}(\textbf{d}+(-1,1,0)),\frac{1}{2}(\textbf{d}+(1,-1,0))\}$.
@@
    For each $r\geq 2$, $\mathcal{W}=\{(-1^{2}),\dots,(-1^{r}),(1^{1},-2^{2}),\dots,(1^{1},-2^{r})\}$, with $d_{1}=1$ and $d_{i}=2m_{i}$ for positive integers $m_{i}$, $i=2,\dots,r$. This is the Bérard-Bergery–Calabi ansatz, discussed in [2, Example 18]. We have a superpotential with
    $\mathcal{C}=\{\frac{1}{2}(\textbf{d}+(-1^{1})),\frac{1}{2}(\textbf{d}+(1^{1},-2^{2})),\dots,\frac{1}{2}(\textbf{d}+(1^{1},-2^{r})),\frac{1}{2}(\textbf{d}+(1^{1}))\}$.
@@
Remark. As in the Ricci-flat case, the only superpotential for $r>3$ is for the Bérard-Bergery–Calabi ansatz. For $r=2,3$, only cases (2) and (3) of Theorem 1.14 of [6] are allowed, with further restrictions on the dimension.

### v2

@@ 3.1. Classification of superpotentials
together imply that $v_{1}=\pm 1$, $v_{2}=\pm 1$, $d_{1}=d_{2}=2$, and also that $\textbf{v}=\pm(1^{1},-1^{2})$. This implies that all vertices of $\mathrm{conv}(\mathcal{W})$ lie in the two planes
@@
Thus, there can only be type III vertices in $\mathrm{conv}(\mathcal{W})$. In fact, there are no type II vectors in $\mathcal{W}$. To see this, assume for sake of contradiction that $(1^{1},-1^{2},-1^{3})\in\mathcal{W}$. Then, by Proposition 4.1(i) and (vii) of [4], $(1^{1},-2^{2}),(-2^{1},1^{2}),(1^{1},-2^{3}),(-2^{1},1^{3}),(1^{2},-2^{3}),\\ (-2^{2},1^{3})\in\mathcal{W}$. As a result, $d_{1}=d_{2}=d_{3}=4$, and (3.4),(3.5) require that for each $i,j=1,2,3$ with $i\neq j$,
@@
On the other side, we first suppose that $(1^{1},-1^{2},-1^{i})$ is a vertex. Then $d_{1}=2$, in which case all vertices $\textbf{w}^{\prime}$ of $\mathrm{conv}(\mathcal{W})$, with the possible exception of $\textbf{w}_{0}$, must satisfy $\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}=1$. However, this is impossible, as $(-1^{2})\in\{\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}<1\}$ but $\textbf{w}_{0}\in\{\frac{w^{\prime}_{1}}{2}-\frac{w^{\prime}_{2}}{2}>1\}$. The only remaining possibility is that $(1^{1},-2^{i})$ is a vertex, so that $d_{1}=1$ by (3.8). Similarly, $(-1^{2})\in\{w^{\prime}_{1}-\frac{w^{\prime}_{2}}{2}>1\}$ and $(1^{1},-2^{2})\in\{w^{\prime}_{1}-\frac{w^{\prime}_{2}}{2}<1\}$, which is a contradiction. Therefore there are no type II vectors in $\mathcal{W}$ with nonzero entries in the first two positions.
Consider the plane spanned by the first 3 coordinates (see Figure 1). There are no type II vectors, and furthermore (3.8) implies that $(-2^{1},1^{3}),(-2^{1},1^{2})\notin\mathcal{W}$. Since $(-1^{1})$ cannot be a vertex, we have $d_{1}=1$. Let $\textbf{c}_{\pm}=\frac{1}{2}(\textbf{d}+\textbf{v}_{\pm})$. Since $\textbf{v}_{+}+\textbf{v}_{-}=2\textbf{w}_{0}=(2^{1},-4^{2})$, we must in particular have $v_{+,1}+v_{-,1}=2$, but also since both $\frac{1}{2}(\textbf{d}+\textbf{v}_{\pm})$ are null and $d_{1}=1$, we must have $v_{\pm,1}^{2}<1$, which is a contradiction. Thus $(a,b)=(1,1)$ is impossible. ∎
@@
If $(a,b)=(1,0)$, then the only set of weight vectors admitting a superpotential is the Bérard Bergery–Calabi ansatz.
@@
Let $\frac{1}{2}(\textbf{d}+\textbf{w})$ be a (non-null) vertex adjacent to $\frac{1}{2}(\textbf{d}+\textbf{v})$. It is also adjacent to $\frac{1}{2}(\textbf{d}-\textbf{v})$, so that $J(\textbf{d}-\textbf{v},\textbf{d}+\textbf{w})=0$. This implies that $J(\textbf{d}+\textbf{v},\textbf{d}+\textbf{w})\neq 0$, so either $\frac{1}{2}(\textbf{d}+\textbf{w})$, $\frac{1}{2}(\textbf{d}+\textbf{v})$ do not satisfy the unique sum condition, or $\frac{1}{2}(\textbf{v}+\textbf{w})\in\mathcal{W}$.
@@
Since $\mathrm{conv}(\mathcal{W})$ has nonzero $(r-1)$-dimensional measure, there are at least $(r-1)$ vertices in $\{s=-1\}$ adjacent to $\frac{1}{2}(\textbf{d}+\textbf{v})$, which we denote by $\frac{1}{2}(\textbf{d}+\textbf{w}_{1}),\dots,\frac{1}{2}(\textbf{d}+\textbf{w}_{r-1})$. For each $i$, we would then have that $\textbf{w}^{\prime}_{i}\equiv\frac{1}{2}(\textbf{v}+\textbf{w}_{i})\in\mathcal{W}$. Thus
@@
Thus the $\{\frac{1}{2}(\textbf{d}+\textbf{w}_{i})\}$ and the $\{\frac{1}{2}(\textbf{d}+\textbf{w}^{\prime}_{i})\}$ lie on parallel affine hyperplanes of $\{s=-1\}$.
@@
Consequently, $d_{1}=d_{2}=2$, $v_{1}=-1$, $v_{2}=1$, and the right-hand side inequality is an equality, so $\textbf{v}=(-1^{1},1^{1})$, contradicting $\sum_{i=1}^{r}v_{i}=-1$. Hence, all three of $v_{1},v_{2},v_{3}$ are equal to ±$\pm 1$. In fact, since $d_{1},d_{2},d_{3}\geq 2$, we must have $v_{1}=-1,v_{2}=1,v_{3}=1$, and the null condition inequality is an equality, implying that $\textbf{v}=(-1^{1},1^{2},1^{3})$, contradicting $\sum_{i=1}^{r}v_{i}=-1$.
Hence, by elimination, each $\textbf{w}_{i}$ is a type III vector. Suppose without loss of generality that $\textbf{w}_{1}=(1^{1},-2^{2})$. Then
@@
Suppose for now that $r\geq 3$, and consider the equality
@@
The only way in which the difference of two type III vectors can have only even entries is if the $(-1)$’s are in the same position. Since there must be $r-1$ such vectors, the only possibility up to permutation is $\textbf{w}_{i}=(1^{1},-2^{i+1})$, $i=1,\dots,r-1$. All the $\textbf{w}_{i}$’s lie on the hyperplane $\{x_{1}=1\}\cap\{\sum_{i=1}^{r}x_{i}=-1\}$. Since $2\textbf{w}^{\prime}_{i}-\textbf{w}_{i}=\textbf{v}$ for each $i$, we must have $(\textbf{w}^{\prime}_{1})_{1}=(\textbf{w}^{\prime}_{2})_{1}=\dots=(\textbf{w}^{\prime}_{r-1})_{1}$.
@@
First, $d_{1}=1$. Then $v_{1}\in\{-1,0\}$. Since $v_{1}$ must differ from $-(\textbf{w}_{i})_{1}=-1$ by an even integer, we have $v_{1}=-1$. Since $\frac{1}{2}(\textbf{d}+\textbf{v})$ is null, we have $\textbf{v}=(-1,0,\dots,0)$ and we have exactly Example 18 of [2], on the Bérard Bergery–Calabi ansatz.
@@
In particular, we must have $|v_{1}/d_{1}|\leq 1$, with equality if and only if $v_{1}=-d_{1}$ and $\textbf{v}=(-d_{1},0,\dots,0)$. However, we have assumed that $d_{1}>1$, so equality is impossible. Therefore $|v_{1}/d_{1}|<1$ and $v_{k}/2>0$ for each $k$. The only possibility is $v_{k}=1$ for each $k$, with $v_{1}/d_{1}=-1/2$. The average of $v_{1}$ and $1$ must be a component of a vector in $\mathcal{W}$, so in fact $v_{1}\in\{-1,-3,-5\}$. The second equation implies that in fact $v_{1}=-1$, $d_{1}=2$, so that $\sum_{i=1}^{r}v_{i}\geq 0$, which is a contradiction.
In the case $r=2$, there are only three possibilities for $\textbf{w}^{\prime}_{1}$. First, if $\textbf{w}^{\prime}_{1}=(0,-1)$, then $\textbf{v}=(1,0)$ and we again have the Bérard Bergery–Calabi ansatz. Otherwise, either $\textbf{w}_{1}^{\prime}=(0,-1)$ and $\textbf{v}=(-3,2)$ or $\textbf{w}_{1}^{\prime}=(-2,1)$ and $\textbf{v}=(-5,4)$; in both of these cases, the null condition for $\frac{1}{2}(\textbf{d}+\textbf{v})$ contradicts (3.10). This concludes the classification for the case $(a,b)=(1,0)$. ∎
@@
Let the two (adjacent) vertices in $\{s>-1\}$ be $\frac{1}{2}(\textbf{d}\pm\textbf{v})$ and let $\frac{1}{2}(\textbf{d}+\textbf{u})$ be the vertex in $\{s<-1\}$. Let $Q_{\pm}=\ker J(\textbf{d}\pm\textbf{v},\cdot)\cap\{s=-1\}$ and $Q_{u}=\ker J(\textbf{d}+\textbf{u},\cdot)\cap\{s=-1\}$. Each is a hyperplane of $\{s=-1\}$, and $Q_{+}\cap Q_{-}=\emptyset$.
@@
We consider the situation case by case, choosing each of $\textbf{w}_{\pm}$ to be a type I, II or III vector in turn.
@@
Thus, we have obtained a special case of the Bérard Bergery–Calabi ansatz discussed in [2]. We permute the indices $1\leftrightarrow 3$ to match the notation, so that the geometric data is $(d_{1},d_{2},d_{3})=(1,2,2)$ and $\mathcal{W}=\{(0,-1,0),(0,0,-1),(1,-2,0),(1,0,-2)\}$. However, the superpotential vectors differ from those in Lemma 3.15 and [2, Example 18]. They are $\mathcal{C}=\frac{1}{2}(\textbf{d}+(\mathcal{B},0))$, where
@@
Suppose $\textbf{w}_{+}$ is of type II, say $\textbf{w}_{+}=(1^{1},-1^{2},-1^{3})$, and that $\textbf{w}_{-}$ is also of type II and thus lies in $R$. Say $\textbf{w}_{-}=(-1^{1},1^{2},-1^{3})$, then $\textbf{u}=(-2^{3})$ and $\textbf{v}=(2^{1},-2^{2})$. The null condition requires that $d_{3}=4$. The conditions $J(\textbf{d}+\textbf{w}_{+},\textbf{d}+\textbf{w}_{-})=0$ and $J(\textbf{d}\pm\textbf{v},\textbf{d}\pm\textbf{v})=0$ are equivalent and both reduce to $\frac{4}{d_{1}}+\frac{4}{d_{2}}=1$. Consider again the plane spanned by the first three coordinates. Since $\overline{\textbf{w}_{+}\textbf{w}_{-}}$ is an edge, we have $(1^{1},-2^{2}),(1^{1},-2^{3})\notin\mathcal{W}$. Furthermore, $(1^{1},-2^{2},0)$, $(0,-2,1)$ and $(-1^{1},-1^{2},1^{3})$ cannot be vertices adjacent to $\textbf{w}_{+}$, since they are not in $\ker J(\textbf{d}+\textbf{w}_{+},\textbf{d}+\cdot)$. This is a contradiction, as can be seen from Figure 1 and the fact that $(-1^{1},-1^{2},1^{3})\in\mathcal{W}$.
@@
Since $(-1^{2})\in\mathcal{W}$ by Lemma 4.1(ii) of [4], and we have eliminated all other possibilities as can be seen in Figure 1, $(-1^{2})$ must be adjacent to $\textbf{w}_{+}$, and therefore $1=\frac{2}{d_{2}}$ and $d_{2}=2$, contradicting (3.14) and completing the proof of the lemma. ∎
We may summarise Lemmas 3.13, 3.14, 3.15 and 3.16 in the following statement. Superpotentials (1)-(4) had already been found in [2].
@@
    $\mathcal{W}=\emptyset$ with $d_{1}=1$. This is the 2-dimensional Bryant soliton, discussed in [2, Example 15]. There is one superpotential, with $\mathcal{C}=\frac{1}{2}(\textbf{d}+\{(1,0),(-1,0)\})$.
@@
    $\mathcal{W}=\{(-1)\}$ with $d_{1}=4$. This is the 5-dimensional Bryant soliton, discussed in [2, Example 15]. There is one superpotential, with $\mathcal{C}=\frac{1}{2}(\textbf{d}+\{(0,0),(-2,0)\})$.
@@
    $\mathcal{W}=\{(-1,0),(0,-1)\}$ with $d_{1}=d_{2}=2$. This is the warped product of two Einstein 2-manifolds with positive Einstein constant, discussed in [2, Example 16]. There is one superpotential, with $\mathcal{C}=\frac{1}{2}(\textbf{d}+\{(-1,-1,0),(-1,1,0),(1,-1,0)\})$.
@@
    For each $r\geq 2$, $\mathcal{W}=\{(-1^{2}),\dots,(-1^{r}),(1^{1},-2^{2}),\dots,(1^{1},-2^{r})\}$, with $d_{1}=1$ and $d_{i}=2m_{i}$ for positive integers $m_{i}$, $i=2,\dots,r$. This is the Bérard Bergery–Calabi ansatz, discussed in [2, Example 18]. We have a superpotential with $\mathcal{C}=\frac{1}{2}(\textbf{d}+\{(-1)^{1},(1^{1},-2^{2}),\dots,(1^{1},-2^{r}),(1^{1})\})$.
@@
Remark. As in the Ricci-flat case, the only superpotential for $r>3$ is for the Bérard Bergery–Calabi ansatz. For $r=2,3$, only cases (2) and (3) of Theorem 1.14 of [6] are allowed, with further restrictions on the dimension.